OpenStudy (anonymous):

solve the radical equation... (x+15)1/4=(2x)1/4

OpenStudy (anonymous):

multiply the perenthesis by 1

OpenStudy (anonymous):

so the answer is (15,1)

OpenStudy (anonymous):

wait wat?

OpenStudy (anonymous):

so.....

OpenStudy (anonymous):

how did u get (15,1)

OpenStudy (anonymous):

we have (1/4) on both sides of the equation, we can just multiply both sides by 4. we will be left with (x+15)=2x ===>combine like turns==> subtract x from both sides ==> 15=2x-x ==> x=15

OpenStudy (anonymous):

ohh eq editor ...\[(x+15)(1/4)=(2x)(1/4)\] ==> \[(x+15)=2x\] ==> \[15=2x-x\] ==> \[x=15\]